The velocity of a diver just before hitting the water is -11.7 m/s, where the minus sign indicates that her motion is directly downward. What is her displacement during the last 1.16 s of the dive?
All steps please!
Answer #1

As mentioned in the problem,

Initial velocity of the diver, u = -11.7 m/s

g = gravitational acceleration = -9.81 m/s^2

t = 1.16 s

Expression for the displacement is given as,

d = u*t + (1/2)*g*t^2

=> d = -11.7*1.16 + (0.5)(-9.81)(1.16^2)

= -13.57 - 6.60

= -20.17 m

Hence displacement of the diver will be d = -20.17 m at last 1.16 s.

Negative sign shows that the location will be inside the water.